Determine the radius of the spherical wedge whose volume 12 m3 with a central angle of 1.80 radians.


1
Expert's answer
2021-09-26T11:59:20-0400

"V_{wedge} = \frac{2}{3} R^3 \theta_{rad}"


"V_{wedge}=12m^3"

"\theta_{rad}=1.80"

so we will get


"\frac{12\times3/2}{1.80}=R^3"


"R=2.15443\space answer"